OpenStudy (anonymous):

The value of Mr. Dulaney's car x years after its purchase is given by the function, V(x) = 15,000(0.87)^x. Approximately, what was the value of Mr. Dulaney's car 5 years after its purchase? a. $7, 500 b. $8, 600 c. $9, 900 d. $13, 100

OpenStudy (anonymous):

V(x)=15,000*(0.87)^x V(5)=15,000*(0.87)^5 V(5)=15,000*(~0.4984).... that approximatly not a negative just so we are clear V(5)=7,476.31 so i would say a.